ABC Corporation has three service departments with the following costs and activity base:

Service Department Cost Activity Base for Allocation
Graphics Production $200,000 number of copies
Accounting 500,000 number of invoices processed
Personnel Department 400,000 number of employees


ABC has three operating divisions, Micro, Macro and Super. Their revenue, cost and activity information are as follows:

Micro Macro Super
Direct revenues $700,000 $850,000 $650,000
Direct operating expenses 50,000 70,000 100,000
Number of copies made 20,000 30,000 50,000
Number of invoices processed 700 800 500
Number of employees 130 145 125


How much service department cost will be allocated to the Micro Division?

a.$145,000

b.$60,000

c.$345,000

d.$200,000
Answer #1

OPTION d: 345000

EXPLANATION:

Total service department cost to be allocated to Micro Division:

Particulars Amount
Graphic production cost - based on number of copies made $200000 x (20000/100000) = $40000
Accounting cost - baseed on number of invoices processed $500000 x (700/2000) = $175000
Personnel Department cost - based on number of employees $400000 x (130/400) = $130000
Total $345000